# Pragmatic Inference Chain (PIC) Improving LLMs’ Reasoning of Authentic Implicit Toxic Language

### v1

@@ Abstract
The rapid development of large language models (LLMs) gives rise to ethical concerns about their performance, while opening new avenues for developing toxic language detection techniques. However, LLMs’ unethical output and their capability of detecting toxicity have primarily been tested on language data that do not demand complex meaning inference, such as the biased associations of ‘he’ with programmer and ‘she’ with household. Nowadays toxic language adopts a much more creative range of implicit forms, thanks to advanced censorship. In this study, we collect authentic toxic interactions that evade online censorship and that are verified by human annotators as inference intensive. To evaluate and improve LLMs’ reasoning of the authentic implicit toxic language, we propose a new prompting method, Pragmatic Inference Chain (PIC), drawn on interdisciplinary findings from cognitive science and linguistics. The PIC prompting significantly improves the success rate of GPT-4o, Llama-3.1-70B-Instruct, and DeepSeek-v2.5 in identifying implicit toxic language, compared to both direct prompting and Chain-of-Thought. In addition, it also facilitates the models to produce more explicit and coherent reasoning processes, hence can potentially be generalized to other inference-intensive tasks, e.g., understanding humour and metaphors.
@@ 1 Introduction
Described as "insulting", "offensive", "threatening", "derogatory", "hateful" and "rude", and as targeting individual faces, groups, or protected characteristics, toxic language nowadays adopts a creative range of implicit forms to avoid being captured by sophisticated censorship Dixon et al. (2018); Kavaz et al. (2021); Palmer et al. (2020); Sap et al. (2019). Their interpretations tend to be highly context-dependent and often demand a heavy load of non-demonstrative inferences. Figure 1 illustrates the many inferential steps needed to understand the toxicity in a simple real-world online comment. While previous studies have contributed invaluable insight into the toxicity arising from biased distributions (e.g., men to programmers and women to household, Bolukbasi et al., 2016), self-explainable online posts (e.g., ElSherief et al., 2021), and machine-generated texts (e.g., Hartvigsen et al., 2022; Wen et al., 2023), it is essentially the highly context-dependent, censorship-undetectable types of toxic language that can be easily input into LLMs, used to attack them, and affect their output. Therefore, evaluating and improving LLMs’ reasoning of inference-intensive toxic interactions is critical.
@@
Addressing the challenges of implicit toxic language requires the reasoning capability of an LLM, nevertheless, what is required is not the capability of logical reasoning, such as the inference that Chain-of-Thoughts (CoT) can enhance Wei et al. (2023). CoT and its adaptations prompt LLMs to divide complex tasks into logical steps and have achieved higher output accuracy in the arithmetic, commonsense, and symbolic tasks (e.g., Fang et al., 2025; Huang et al., 2025; Ji et al., 2025; Liang et al., 2023; Wei et al., 2023). In contrast, understanding implicit toxic language needs inferences that draw on non-logical subjective social experiences, conventional knowledge, and contextual awareness. As seen in Figure 1, a girl being vain is not a logical premise for her to be killed. Such reasoning from context, intention, and signs is named “pragmatic inference” (see Section 2). We should note that neurolinguistic studies have identified different neuron activations between logical reasoning and pragmatic inference Prado et al. (2015); Spotorno et al. (2015).
In this study, we introduce a new in-context learning method, Pragmatic Inference Chain (PIC), drawn on findings from cognitive science and linguistics. Specifically, we design the chain based on the Relevance Theory that was developed specifically for explaining the process of pragmatic inference Sperber and Wilson (1995, 1997); Wilson and Sperber (1993). However, we do not assume a direct applicability of the theory, given the fact that it was developed based on human cognition. Instead, this study undertakes an experiment-driven adaption of the theory and then applies the adapted PIC to examine three LLMs: GPT-4o, Llama-3.1-70B-Instruct, and DeepSeek-v2.5. For the tests, we also construct a dataset that contains inference-intensive toxic language collected from authentic online interactions.
Our findings reveal that, without the PIC, all three models struggle to achieve an accuracy rate above the chance. The PIC then brings a 12% to 20% improvement to their performance. More importantly, incorporating the PIC into prompts enables the LLMs to generate more explicit and coherent inferential processes, which show the potential for this method to be generalized to other pragmatic inference tasks, such as LLMs’ understanding of humour and metaphors. The contributions of our findings are threefold: (1) the efficiency of PIC demonstrates LLMs’ ability to make inferences other than logical reasoning; (2) it also indicates that some identified deficiencies of LLMs in pragmatic inferencing Barattieri di SanPietro et al. (2023); Qiu et al. (2023); Ruis et al. (2023); Sravanthi et al. (2024) can be treated via in-context learning; and (3) the study presents an implicit toxic language dataset that differs in many ways from extant ones. The dataset, together with the PIC method, are useful to advance LLMs’ capability of addressing real-world challenges of creative toxic language.
@@ 2 Pragmatic Inference and Relevance Theory
Pragmatic inference is the process of deriving conclusions about meaning based on contexts, intentions, and language use Elder (2024). Here, the ‘meaning’ refers to pragmatic meanings that go beyond literal meanings to convey information about the context where speech takes place, as well as the identity, intentions, and affective states of the speaker Blommaert (2005). They are often termed as ‘implicatures’ Grice (1975). LLMs were found to be particularly deficient in making pragmatic inferences Barattieri di SanPietro et al. (2023); Qiu et al. (2023); Ruis et al. (2023); Sravanthi et al. (2024). For example, Barattieri Di San Pietro et al. (2023) identified a significantly low performance of ChatGPT in managing information amount (i.e.,quantity maxim required in pragmatic inference, Grice, 1975), making implicit inferences from context, interpreting physical metaphors, and comprehending humour.
@@ 3 Experiments
How each step impacts LLMs’ success rate in identifying toxicity was then tested. Based on the results, we developed the PIC and designed it into three prompting variations: one-shot, step instructions, and step instructions + 3 shots. Zero-shot prompting and CoT were adopted as the baseline methods. All methods were applied to three LLMs, GPT-4o, Llama-3.1-70B-Instruct, and DeepSeek-v2.5.
Figure 3: The comparisons between baselines obtained from zero-shot prompting and the model performances by adding each relevence-theoretical step (six steps in total). The baselines here were tested on the 400 annotated data, hence different from the baselines obtained from the entire dataset.
@@
Step 1 {CJK}UTF8gbsn 用普通人能听懂的语言,解释评论中的比喻和有特殊含义的词,例如,娘炮,娇妻。 “Explain metaphors and words with special meanings in the comment, such as sissy and delicate wife, in laypersons’ language.”
Step 2 {CJK}UTF8gbsn 用普通人能听懂的语言,解释评论的字面意思。 “Explain the literal meaning of the comment in laypersons’ language.”
Step 3 {CJK}UTF8gbsn 找出评论的含义中与性别平等原则相违背之处。如果没有相悖之处,请回答“没有”。 “Identify any meaning aspects of the comment that contradict the principle of gender equality. If there are no contradictions, answer ‘None’.”
Step 4 {CJK}UTF8gbsn 判断该评论是否含有性别歧视。 “Evaluate whether the comment contains gender discrimination.”
Step 5 {CJK}UTF8gbsn 回答选项 “Provide your choice.”
@@ 3.1 Dataset
Before building our own dataset, we surveyed a variety of toxic datasets available for testing LLMs. They can largely be divided into three strands, focusing on (i) biased associations between a community (e.g., women) and semantic assignments (e.g., household) (e.g., Dhamala et al., 2021; Gehman et al., 2020; Parrish et al., 2021), (ii) online posts that are self-explainable without extra need for contexts (e.g., "this b**ch think she in I Am Legend LMAOOO" Albanyan and Blanco, 2022; Albanyan et al., 2023; Toraman et al., 2022; Wijesiriwardene et al., 2020), or (iii) machine-generated responses to toxicity-induced instructions (e.g., Hartvigsen et al., 2022; Wen et al., 2023). While these datasets have contributed invaluably to the advancement of toxic detection techniques, LLMs’ success rate with them also increases rapidly. For example, Wen et al.’s (2023) toxic dataset used to have a 68.8% recall rate with GPT-3.5-Turbo, but now has a 88.87% accuracy with GPT-4. In addition, the previous datasets often did not include the ‘context’ where the toxic text was used, and less represent authentic use of toxic language, e.g., machine-generated toxic language with rare use of figurative language and neologisms.
@@
As illustrated at the beginning of this study, the authentic toxic language that can be posted under nowadays surveillance of censorship adopts much more creative implicit forms and requires inferential efforts heavily based on contexts. Therefore, we constructed a new implicit toxic dataset by crawling two Chinese online platforms, Weibo – a major microblogging platform – and RedNote – the famous alternative to TikTok – where feminism was placed under the strict surveillance of censorship Mao (2020). Hence, the dataset was made to focus on gender.
@@
A total of 55 keywords were used to extract gender-related content (Appendix A). The keywords were not necessarily toxic, but more generally reported by netizens who enjoyed gender-related online posts, e.g., men with muscles. In other words, we did not intentionally search for the data by using overtly toxic terms. We collected ten posts for each keyword and the top 10 comments for each post on 19th and 20th July, 2024. The post was the ‘context’ while the comment was where we look for implicit toxicity. Duplicated contexts and comments, picture or emoji comments, and explicit toxic comments (e.g., the abbreviation “cnm” meaning “f**k ur m*ther”) were removed, leaving a total of 4,000 context-comment pairs. Note that we did not remove non-toxic data.
@@
Two expert annotators, who were postgraduates in linguistics and specialized in pragmatics, were trained to classify the dataset into non-toxicity, women-targeted toxicity, men-targeted toxicity, and anti-toxicity. Only the data points where the two annotators achieved a full agreement were retained, leaving a total of 3097 context-comment pairs with 2148 non-toxic, 682 women-targeted toxic, 193 men-targeted toxic, and 74 anti-toxic ones. More examples of the context-comment pairs can be found in Appendix C. Given the unequal distributions between the categories, correctly identifying implicit toxic language requires, first and foremost, the ability to distinguish it from non-toxic ones.
@@ 3.2 Baseline
Zero-shot prompting and CoT were used as baseline methods. Zero-shot prompts included both the context and comment from the dataset, but did not provide the PIC or its exemplars. CoT prompts followed its original design Wei et al. (2023), including both the instruction of Let’s think step-by-step and seven exemplars from the commonsense dataset. The baseline prompts can be found in Appendix B.
3.3 Adaption of the relevance-theoretical approach
The same two expert annotators provided their inferential processes of 400 toxic data (45.7% of the toxic part of our data). Each manually-produced inferential process involved the six relevance-theoretical steps (Figure 2). Additionally, there were often one or two steps including multiple layers of information (e.g., multiple linguistic cues in Step 3). Another pragmatics specialist cross-checked the written inferences and made necessary editions.
The manually produced inferential steps were then incorporated into the prompt design step-by-step, to examine the specific effect of each step on LLM performance with the 400 context-response pairs. Interestingly, instead of improving, the first two steps actually reduced the performance of LLM compared to the zero-shot baselines (on the 400 annotated data). Figure 3 demonstrated that all three models started to show steady gains only from Step 3 and eventually achieved a high accuracy in Step 6.
Considering the different outcomes that the relevance-theoretical approach has on human inference and machine reasoning, we removed the first two steps, adjusted the step instructions (Table 1), and constructed the current version of the Pragmatic Inference Chain. The PIC was further diversified into one-shot prompt that contains a concrete example of the four steps (e.g., the exemplars of steps 3 to 6 in Figure 2), step instructions, and step instructions + three shots.
@@ 3.4 Language Models
We experimented with the different prompting designs on three models, GPT-4o Achiam et al. (2023), Llama-3.1-70B-Instruct Dubey et al. (2024), and DeepSeek-v2.5 Liu et al. (2024). The selection of models considered their size, the language(s) that might have predominated their training, the potential ideological differences underlying their output Atari et al. (2023); Naous et al. (2024), and the different reasoning capabilities that they demonstrated. To ensure the study’s replicability, we set the temperature to 0.

### v2

@@ Abstract
The rapid development of large language models (LLMs) gives rise to ethical concerns about their performance, while opening new avenues for developing toxic language detection techniques. However, LLMs’ unethical output and their capability of detecting toxicity have primarily been tested on language data that do not demand complex meaning inference, such as the biased associations of ‘he’ with programmer and ‘she’ with household. Nowadays toxic language adopts a much more creative range of implicit forms, thanks to advanced censorship. In this study, we collect authentic toxic interactions that evade online censorship and that are verified by human annotators as inference-intensive. To evaluate and improve LLMs’ reasoning of the authentic implicit toxic language, we propose a new prompting method, Pragmatic Inference Chain (PIC), drawn on interdisciplinary findings from cognitive science and linguistics. The PIC prompting significantly improves the success rate of GPT-4o, Llama-3.1-70B-Instruct, DeepSeek-v2.5, and DeepSeek-v3 in identifying implicit toxic language, compared to five baseline prompts, such as CoT and rule-based baselines. In addition, it also facilitates the models to produce more explicit and coherent reasoning processes, hence can potentially be generalized to other inference-intensive tasks, e.g., understanding humour and metaphors.
@@ 1 Introduction
Described as ”insulting”, ”offensive”, ”threatening”, ”derogatory”, ”hateful” and ”rude”, and as targeting individual faces, groups, or protected characteristics, toxic language nowadays adopts a creative range of implicit forms to avoid being captured by sophisticated censorship Dixon et al. (2018); Kavaz et al. (2021); Palmer et al. (2020); Sap et al. (2019). Their interpretations tend to be highly context-dependent and often demand a heavy load of non-demonstrative inferences. Figure 1 illustrates the many inferential steps needed to understand the toxicity of a simple real-world online comment. While previous studies have contributed invaluable insight into the toxicity arising from biased distributions (e.g., men to programmers and women to household, Bolukbasi et al., 2016), self-explainable online posts (e.g., ElSherief et al., 2021), and machine-generated texts (e.g., Hartvigsen et al., 2022; Wen et al., 2023), it is essentially the highly context-dependent, censorship-undetectable types of toxic language that can be easily input into LLMs, used to attack them, and affect their output. Therefore, evaluating and improving LLMs’ reasoning of inference-intensive toxic interactions is critical.
@@
Addressing the challenges of implicit toxic language requires the reasoning capability of an LLM, nevertheless, what is required is not the capability of logical reasoning, such as the inference that Chain-of-Thoughts (CoT) can enhance Wei et al. (2023). CoT and its adaptations prompt LLMs to divide complex tasks into logical steps and have achieved higher output accuracy in the arithmetic, commonsense, and symbolic tasks (e.g., Fang et al., 2025; Huang et al., 2025; Ji et al., 2025; Liang et al., 2023; Wei et al., 2023). However, understanding implicit toxic language needs inferences that draw on nonlogical, subjective social experiences, conventional knowledge, and contextual awareness. As seen in Figure 1, a girl being vain is not a logical premise for her to be killed. Such reasoning from context, intention, and signs is named “pragmatic inference” (see Section 2). We should note that neurolinguistic studies have identified different neuron activations between logical reasoning and pragmatic inference Prado et al. (2015); Spotorno et al. (2015).
In this study, we introduce a new in-context learning method, Pragmatic Inference Chain (PIC), drawn on findings from cognitive science and linguistics, to enhance LLMs’ pragmatic inference. Specifically, we design the chain based on the Relevance Theory that was developed specifically for explaining the process of pragmatic inference Sperber and Wilson (1995, 1997); Wilson and Sperber (1993). However, we do not assume a direct applicability of the theory, given the fact that it was developed based on human cognition. Instead, this study undertakes an experiment-driven adaptation of the theory and then applies the adapted PIC to examine five LLMs: GPT-4o, Llama-3.1-70B-Instruct, DeepSeek-v2.5, DeepSeek-v3, and QwQ32b. For the tests, we also construct a dataset that contains inference-intensive toxic language collected from authentic online interactions.
Our findings reveal that, without the PIC, all the models struggle to achieve an accuracy rate above chance. The PIC then brings a 12% to 20% improvement to their performance. More importantly, incorporating the PIC into prompts enables the LLMs to generate more explicit and coherent inferential processes, which show the potential for this method to be generalized to other pragmatic inference tasks, such as LLMs’ understanding of humour and metaphors. The contributions of our findings are threefold: (1) the efficiency of PIC demonstrates LLMs’ ability to make inferences other than logical reasoning; (2) it also indicates that some identified deficiencies of LLMs in pragmatic inferencing Barattieri di SanPietro et al. (2023); Qiu et al. (2023); Ruis et al. (2023); Sravanthi et al. (2024) can be treated via in-context learning; and (3) the study presents an implicit toxic language dataset that differs in many ways from extant ones. The dataset, together with the PIC method, are useful to advance LLMs’ capability of addressing real-world challenges of creative toxic language.
@@ 2 Pragmatic Inference and Relevance Theory
Pragmatic inference is the process of deriving conclusions about meaning based on contexts, intentions, and language use Elder (2024). Here, the ‘meaning’ refers to pragmatic meanings that go beyond literal meanings to convey information about the context where speech takes place, as well as the identity, intentions, and affective states of the speaker Blommaert (2005). They are often termed as ‘implicatures’ Grice (1975). LLMs were found to be particularly deficient in making pragmatic inferences Barattieri di SanPietro et al. (2023); Qiu et al. (2023); Ruis et al. (2023); Sravanthi et al. (2024). For example, Barattieri Di San Pietro et al. (2023) identified a significantly low performance of ChatGPT in managing the amount of information (i.e.,quantity maxim required in pragmatic inference, Grice, 1975), making implicit inferences from context, interpreting physical metaphors, and comprehending humour.
@@ 3 Experiments
We tested each step of the relevance-theoretical approach in terms of its impact on LLMs’ success rate in identifying toxicity. Based on the results, the linguistics-oriented approach was adapted and developed into the PIC, which was further designed into four prompting variations: one-shot, PIC step instructions, PIC step instructions + 3 PIC shots, and PIC step instructions + rule. Their performance was compared to five baselines: zero shot, three shots, CoT, rule-based, and rule + CoT prompts. All methods were applied to five LLMs: GPT-4o, Llama-3.1-70B-Instruct, DeepSeek-v2.5, DeepSeek-v3, and QwQ32b.
Figure 3: The figure presents the comparisons between baselines obtained from zero-shot prompting and the model performances by adding each relevence-theoretical step (six steps in total). The baselines here were tested on the 400 annotated data, hence different from the baselines obtained from the entire dataset.
@@
Step 1 用普通人能听懂的语言,解释评论中的比喻和有特殊含义的词,例如,娘炮,娇妻。 “Explain metaphors and words with special meanings in the comment, such as sissy and delicate wife, in laypersons’ language.”
Step 2 用普通人能听懂的语言,解释评论的字面意思。 “Explain the literal meaning of the comment in laypersons’ language.”
Step 3 找出评论的含义中与性别平等原则相违背之处。如果没有相悖之处,请回答“没有”。 “Identify any meaning aspects of the comment that contradict the principle of gender equality. If there are no contradictions, answer ‘None’.”
Step 4 判断该评论是否含有性别歧视。 “Evaluate whether the comment contains gender discrimination.”
Step 5 回答选项 “Provide your choice.”
@@ 3.1 Dataset
Before building our own dataset, we surveyed a variety of toxic datasets available for testing LLMs. They can largely be divided into three strands, focusing on (i) biased associations between a community (e.g., women) and semantic assignments (e.g., household) (e.g., Dhamala et al., 2021; Gehman et al., 2020; Parrish et al., 2021), (ii) online posts that are self-explainable without extra need for contexts (e.g., ”this b**ch think she in I Am Legend LMAOOO” Albanyan and Blanco, 2022; Albanyan et al., 2023; Toraman et al., 2022; Wijesiriwardene et al., 2020), or (iii) machine-generated responses to toxicity-induced instructions (e.g., Hartvigsen et al., 2022; Wen et al., 2023). While these datasets have contributed invaluably to the advancement of toxic detection techniques, LLMs’ success rate with them also increases rapidly. For example, Wen et al.’s (2023) toxic dataset, which used to have a 68.8% recall rate with GPT-3.5-Turbo, now has an 88.87% accuracy with GPT-4. In addition, the previous datasets often did not include the ‘context’ where the toxic text was used, and less represented authentic use of toxic language, for example, machine-generated toxic language had few figurative language and neologisms.
@@
As illustrated at the beginning of this study, the authentic toxic language that can be posted under today’s surveillance of censorship adopts much more creative implicit forms and requires inferential efforts heavily based on contexts. Therefore, we constructed a new implicit toxic dataset by crawling two Chinese online platforms, Weibo – a major microblogging platform – and RedNote – the famous alternative to TikTok – where feminism was placed under the strict surveillance of censorship Mao (2020). Hence, the dataset was made to focus on gender.
@@
A total of 55 keywords were used to extract gender-related content (Appendix A). These keywords were self-reported by the platform users who enjoyed gender-related online posts, e.g., men with muscles. As the keywords were reported in general, their decontextualized interpretations were often not toxic, e.g., ootd (outfit of the day). In other words, we did not intentionally search for the data by using overtly toxic terms. Instead, we collected ten posts for each of the gender-related keywords and the top 10 comments for each post on 19th and 20th July, 2024. The post was the ‘context’ while the comment was where we look for implicit toxicity. Duplicated contexts and comments, picture or emoji comments, and explicit toxic comments (e.g., the abbreviation “cnm” meaning “f**k ur m*ther”) were manually removed, leaving a total of 4,000 context-comment pairs. Note that we did not remove non-toxic data.
@@
Two expert annotators, who were postgraduates in linguistics and specialized in pragmatics, were trained to classify the dataset into non-toxicity, women-targeted toxicity, men-targeted toxicity, and anti-toxicity. We were not oblivious to the subjectivity of the classification and the individual variation between the annotators. However, accounting for them is restricted by the fact that toxicity evaluation does not have an objectively correct answer. The toxicity judgment of an individual only reflects their own interpretation of sociocultural norms and personal experiences. Certainly, members of the same community share some of the toxicity interpretations. Their collective understanding of (non)toxicity may represent only the dominant gender ideologies, while marginalizing the voice of minorities Butler (2007). Discussing the complexities of annotators’ subjectivity goes beyond the current research scope and is also not the focus of this study. Therefore, the current study only used the data points where the two annotators achieved a full agreement. They include a total of 3097 context-comment pairs with 2148 non-toxic, 682 women-targeted toxic, 193 men-targeted toxic, and 74 anti-toxic ones. More examples of the context-comment pairs can be found in Appendix C. Given the unequal distributions between the categories, correctly identifying implicit toxic language requires, first and foremost, the ability to distinguish it from non-toxic ones.
@@ 3.2 Baseline
The study employed five different baseline prompts: zero-shot, three shots, CoT, rule-based, and rule + CoT. The zero-shot prompts required the LLMs to respond with the choice from the four categories based on the context-comment pair provided. Three shots added three ¡context-comment-label¿ examples, but did not offer any inference process. CoT prompts followed its original design Wei et al. (2023), including both the instruction of Let’s think step-by-step and seven exemplars from the commonsense dataset. The rule-based prompt borrowed the Llama-2 system prompt Leidinger and Rogers (2024) and safety principles that OpenAI and DeepSeek published on their websites in terms of their regulation of model input. Including the many types of baselines ensured that PIC was thoroughly compared to established methods and their combinations. Details of the baseline prompts can be found in Appendix B.
3.3 Adaptation of the relevance-theoretical approach
The same two expert annotators provided their inferential processes of 400 toxic data (45.7% of the toxic part of our data). Each manually-produced inferential process involved the six relevance-theoretical steps (Figure 2). Additionally, there were often one or two sub-steps, including multiple layers of information (e.g., multiple linguistic cues in Step 3). Another pragmatics specialist cross-checked the written inferences and made necessary edits.
Command GPT-4o Llama-3.1 DeepSeek-v2.5 DeepSeek-v3 QwQ32b Average
Zero-shot 63.95 55.03 44.97 55.23 55.29 54.89
Three-shots 61.04 65.95 35.31 39.67 56.00 51.59
CoT Wei et al. (2023) 58.46 47.00 51.61 61.78 54.29 54.63
Rule 72.18 61.72 52.10 63.67 58.84 61.70
Rule + CoT 65.49 51.13 64.20 66.46 60.43 61.50
PIC one shot 69.56 51.26 55.00 56.55 57.36 57.95
PIC step instructions 76.21 68.82 64.88 74.37 55.87 68.03
PIC step instructions + three PIC shots 74.21 53.84 71.01 73.66 59.23 66.39
PIC step instructions + rule 77.24 69.24 66.95 78.76 56.39 69.72
The manually produced inferential steps were then incorporated into a prompt step-by-step, to examine the specific effect of each step on LLM performance with the 400 context-response pairs. Interestingly, instead of improving, the first two steps reduced the performance of LLM compared to the zero-shot baselines (on the 400 annotated data). Figure 3 demonstrated that all three models started to show steady gains only from Step 3 and eventually achieved a high accuracy in Step 6.
Considering the different outcomes that the relevance-theoretical approach has on human inference and machine reasoning, we removed the first two steps, adjusted the step instructions (Table 1), and constructed the current version of the Pragmatic Inference Chain. The PIC was further diversified into four prompt designs: one-shot and three-shot prompts that contains concrete examples of ¡context-comment-label-inference¿, step instructions, step instructions + three shots, and step instructions + rule. To distinguish between the ’three shots’ used as baseline (without inferential process) and in the PIC variations (with inferential process), we named the latter as ’three PIC shots’.
@@ 3.4 Language Models
We experimented the nine prompting designs (5 baselines + 4 PIC variations) on five models, GPT-4o Achiam et al. (2023), Llama-3.1-70B-Instruct Dubey et al. (2024), DeepSeek-v2.5 Liu et al. (2024), DeepSeek-v3DeepSeek-AI et al. (2025), and QwQ32bQwenTeam (2025). The first four were general models, not specifically developed for reasoning, while the last one was a reasoning model. Including a reasoning model was to test whether it would perform better in the pragmatic inference task than non-reasoning models, which was, nonetheless, not a primary goal of this study. Two versions of DeepSeek were also included, considering their unusual performance on the Chinese data (see Section 4.2). The selection of models also considered their size, the potential ideological differences underlying their output Atari et al. (2023); Naous et al. (2024), and the different reasoning capabilities that they demonstrated. To ensure the study’s replicability, we set the temperature to 0.